suppose the market demand supply are given by p=160-2q and p=50+3q. a per unit excise tax of 10$ is levied on the sellers. what is the list social welfare deadweight loss due to the tax?
Answer #1

Equilibrium quantity before tax is calculated by equating the demand and supply

160-2q=50+3q. Solving for q we get q=22

When tax is imposed it creates a wedge between price demanders pay and price suppliers receive.

price paid by demanders Pd=160-2q and price received by suppliers Ps=50+3q. Per unit tax=10

So Pd-Ps=10 or 160-2q-50-3q=10. Solving for q we get the equilibrium quantity q*=20. Pd=120 and Ps=110

Deadweight loss= 0.5*(Pd-Ps)*(q-q*)=0.5*10*2=10